# QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?

### v4

@@ Abstract
Credit assignment has remained a fundamental challenge in multi-agent reinforcement learning (MARL). Previous studies have primarily addressed this issue through value decomposition methods under the centralized training with decentralized execution paradigm, where neural networks are utilized to approximate the nonlinear relationship between individual Q-values and the global Q-value. Although these approaches have achieved considerable success in various benchmark tasks, they still suffer from several limitations, including imprecise attribution of contributions, limited interpretability, and poor scalability in high-dimensional state spaces. To address these challenges, we propose a novel algorithm, QLLM, which facilitates the automatic construction of credit assignment functions using large language models (LLMs). Specifically, the concept of TFCAF is introduced, wherein the credit allocation process is represented as a direct and expressive nonlinear functional formulation. A custom-designed coder-evaluator framework is further employed to guide the generation and verification of executable code by LLMs, significantly mitigating issues such as hallucination and shallow reasoning during inference. Furthermore, an IGM-Gating Mechanism enables QLLM to flexibly enforce or relax the monotonicity constraint depending on task demands, covering both IGM-compliant and non-monotonic scenarios. Extensive experiments conducted on several standard MARL benchmarks demonstrate that the proposed method consistently outperforms existing state-of-the-art baselines. Moreover, QLLM exhibits strong generalization capability and maintains compatibility with a wide range of MARL algorithms that utilize mixing networks, positioning it as a promising and versatile solution for complex multi-agent scenarios. The code is available at https://github.com/zhouyangjiang71-sys/QLLM.
Introduction
@@
Recent advances in cooperative multi-agent reinforcement learning (MARL)(Oroojlooy and Hajinezhad 2023) have led to widespread applications in domains such as assembly lines (Kaven et al. 2024), logistics distribution (Yan et al. 2022), autonomous driving (Zhang et al. 2024b), and swarm robotics (Orr and Dutta 2023). In such scenarios, multiple agents learn to interact with the environment collaboratively to accomplish designated cooperative tasks. Given that agents receive a shared team reward, it is essential to accurately assess the individual contribution of each agent to facilitate effective training. A key challenge in this context is the credit assignment problem, where inaccurate attribution of individual contributions may result in the emergence of so-called ”lazy agents” (Liu et al. 2023).
To address the credit assignment problem in multi-agent systems, a variety of solutions have been proposed in the academic community. Attention-based credit assignment methods (Soydaner 2022) incorporate a learnable mechanism that adaptively allocates credit by computing attention weights over agents or their interactions. By explicitly modeling inter-agent dependencies, these methods enable the network to selectively attend to the most influential agents or environmental factors when estimating individual contributions. This dynamic weighting scheme enhances the expressiveness of the credit assignment process, particularly in heterogeneous or partially observable environments.
@@
Figure 1: A comparison between the traditional value decomposition method using a mixing network (Left) and our proposed novel paradigm QLLM leveraging LLMs (Right). The traditional approach employs a neural network to model the nonlinear relationship between local Q-values and the global Q-value, whereas QLLM capitalizes on the extensive knowledge encoded within LLMs to directly generate a training-free credit assignment function.
Policy gradient approaches (Agarwal et al. 2021) tackle credit assignment from a differentiable optimization perspective. Each agent’s policy is updated based on gradients computed from shared rewards, typically combined with baselines or variance reduction techniques to improve learning stability. These methods can naturally handle continuous action spaces and stochastic policies, where credit assignment is implicitly achieved by estimating each agent’s impact on the long-term return, either through joint trajectory sampling or individualized advantage estimation. In addition, value decomposition methods (Sunehag et al. 2017) have become increasingly popular. These methods address the credit assignment problem by constructing a joint value function that can be decomposed into individual utility functions associated with each agent. The decomposition is designed to ensure that the optimal joint policy can be recovered from the individual components, allowing for centralized training while preserving decentralized execution. These methods rely on specific structural assumptions that maintain the consistency between the global value and its local components, thereby enabling efficient coordination and scalable training.
However, most existing approaches predominantly rely on neural networks, either explicitly or implicitly, to perform credit assignment for each agents. A representative example of this is the mixing network employed in value decomposition methods. These approaches often suffer from high training costs, limited generalization ability, and a lack of sematic interpretability (Linardatos, Papastefanopoulos, and Kotsiantis 2020). With the advancement of large language models (LLMs) (Kasneci et al. 2023), new paradigms have emerged to address many long-standing challenges in MARL. By leveraging extensive pre-trained knowledge (Petroni et al. 2019), superior generalization capacity (Anil et al. 2022), and powerful code generation abilities (Xu et al. 2022a), LLMs have demonstrated potential as versatile auxiliary components in reinforcement learning pipelines. Recent studies (Qu et al. 2025) have shown the effectiveness of LLMs in generating dense, agent-specific rewards based on natural language descriptions of tasks and overall team goals. Empirical results reveal that such LLM-driven reward shaping leads to faster convergence and improved policy performance over existing baselines. Furthermore, the generalization and multitask learning capabilities of LLMs allow agents to quickly adapt to novel tasks with limited training samples, thus improving sample efficiency (Sanh et al. 2021). In tasks involving human interaction, LLMs can generate responses that more closely align with human expectations, enhancing interaction quality (Zhang et al. 2023b). Given the extensive applications of LLMs in the field of reinforcement learning, we propose to leverage them to address the credit assignment problem in multi-agent systems.
In this paper, we introduce QLLM, a novel value decomposition method that eliminates the need for a mixing network. To support its construction, we first propose a general coder-evaluator framework for LLM-based code generation. Within this framework, LLMs act as both a code generator and a code evaluator (Desmond et al. 2024). Upon receiving task and code generation instructions, the code generator produces a non-linear function called Training-Free Credit Assignment Function (TFCAF), which serves as a direct substitute for the mixing network in algorithms such as QMIX, VMIX, and RIIT (Rashid et al. 2020b; Su, Adams, and Beling 2021; Hu et al. 2021). Generated directly by the LLMs, this credit assignment function offers several advantages, including no need for training, immunity to high-dimensional state spaces, and enhanced semantic interpretability. Additionally, a code verification mechanism has been designed to ensure that the generated code runs correctly while mitigating hallucination issues (Martino, Iannelli, and Truong 2023) commonly encountered by LLMs. This framework significantly reduces the probability of errors when LLMs process complex textual prompts and enables the generation of a reliable TFCAF without direct interaction with the environment, thereby allowing it to replace the mixing network. Since LLMs modify only the mixing network in MARL algorithms, the TFCAF is applicable to almost any MARL algorithm that relies on mixing networks. An IGM-Gating Mechanism is further introduced to flexibly switch the monotonicity constraint on or off, adapting to tasks where the Individual-Global-Max principle holds or does not. The main contributions of this study are summarized as follows:
@@
    We propose the coder-evaluator framework for LLM-based code generation, which enables LLMs to generate a reliable credit assignment function under zero-shot prompting and without any human feedback, referred to as TFCAF. This framework is generalizable and can be extended to other MARL scenarios enhanced by LLMs.
@@
    QLLM is proposed by combining TFCAF and value function decomposition while remaining compatible with most MARL algorithms that employ mixing networks for credit assignment. Unlike existing methods, QLLM can accurately infer each agent’s contribution without interacting with the environment.
@@
    Extensive experiments demonstrate that QLLM outperforms existing credit assignment algorithms, particularly in scenarios involving high-dimensional state spaces. Moreover, the integration of LLMs endows the contribution assignment among agents with semantic interpretability.
Related Work
Credit Assignment in MARL
@@
In recent years, the credit assignment problem has emerged as a fundamental challenge in multi-agent systems. And policy-based approaches (Yu et al. 2022; Zhang et al. 2024a) were among the first to introduce credit assignment mechanisms. MADDPG (Lowe et al. 2017) extends the DDPG (Lillicrap et al. 2015) framework to multi-agent settings by incorporating centralized training with decentralized execution. This approach allows each agent to learn policies that account for the actions of other agents during training, thereby addressing the non-stationarity inherent in multi-agent environments. MADDPG utilizes an actor-critic architecture with centralized critics, where the centralized critic has access to global information, enabling more effective learning in mixed cooperative-competitive scenarios. Similarly, COMA (Foerster et al. 2018) leverages a counterfactual baseline to compute each agent’s advantage function, following the same actor-critic framework as MADDPG. MAAC (Iqbal and Sha 2019) further enhances policy-based methods by introducing an attention mechanism into the critic network, leading to more precise credit assignment. However, most of these algorithms are not directly applicable to tasks with shared rewards.
@@
Another widely adopted paradigm is the value decomposition approach (Rashid et al. 2020b; Xu et al. 2022b), which explicitly decomposes the globally shared return. As one of the earliest methods in this category, VDN (Sunehag et al. 2017) models the global value function as a linear summation of individual value functions. While simple and computationally efficient, VDN struggles to capture complex cooperative relationships due to its linear decomposition. QMIX (Rashid et al. 2020b) improves upon VDN by using a monotonic mixing network to non-linearly combine individual value functions, allowing for more expressive representations. However, the assumption of monotonicity between individual and global values may restrict its flexibility in highly complex environments. QTRAN (Son et al. 2019) and QPLEX (Wang et al. 2020) mitigate this limitation by relaxing the monotonicity constraint, enabling the model to learn richer representations, albeit at the cost of increased computational complexity. LICA (Zhou et al. 2020) adopts an alternative approach by introducing an independent credit assignment network that directly derives individual contributions from global information. Unlike methods focused solely on value decomposition, LICA explicitly models individual contributions, leading to more accurate credit assignment in cooperative tasks. RIIT (Hu et al. 2021) goes further by modeling implicit influence relationships between agents, enabling better credit assignment in tasks with long-term dependencies. Nevertheless, both RIIT and LICA continue to employ a mixing network structure analogous to that used in QMIX, in order to establish the mapping from individual agent Q-values to the global Q-value. Although the mixing network is a critical component for implementing credit assignment, it suffers from limited interpretability and slow convergence during training. Despite the advantages of value decomposition methods, achieving efficient and precise credit assignment in multi-agent systems remains a challenging open research problem.
LLM-enhanced Reinforcement Learning
Leveraging their large-scale pretraining and strong generalization capabilities, LLMs have been increasingly integrated into Reinforcement Learning to enhance performance in multitask learning, sample efficiency, and high-level task planning (Zhang et al. 2023a; Li et al. 2024). In the SMAC-R1 framework (Deng et al. 2024), agents utilize LLMs to generate decision tree code by providing task descriptions. The model further improves the quality and interpretability of the decision tree through self-reflection, guided by feedback from environment-provided rewards. Additionally, LLMs can serve as generators to create high-fidelity multimodal world models, reducing the cost of learning in real-world environments. They can also generate natural language rationales to explain agent behaviors, improving the transparency of decision-making processes. For instance, the Reflexion framework (Shinn et al. 2023) introduces a reinforcement mechanism guided by natural language feedback, without requiring updates to the model parameters. Agents reflect on task-related feedback signals through language and store them in contextual memory, enabling improved decision-making in future interactions. Moreover, LLMs can explicitly or implicitly shape reward signals for RL tasks. The EUREKA framework (Ma et al. 2023) achieves superhuman-level reward function coding through evolutionary search and policy feedback, empowering agents to execute more complex behaviors.
@@
These examples demonstrate the powerful information processing and code generation capabilities of LLMs. Therefore, some researchers have leveraged LLMs to tackle the challenges of temporal and agent-level credit assignment in MARL. Qu et al. (2025) were the first to introduce LLMs into the credit assignment problem. They proposed the notion of latent rewards, making LLMs to capture subtle cooperative relationships among agents and to decompose global rewards accordingly. This method effectively addresses the challenge of multi-agent credit assignment under sparse rewards and outperforms existing baseline algorithms. However, it suffers from instability in code generation and may produce latent rewards that contain redundant or difficult-to-interpret components. Additionally, Lin et al. (2025) employed LLMs to learn potential-based reward functions across multiple queries, allowing for a more accurate estimation of individual agent contributions. These methods primarily focus on decomposing the global reward across agents but remain limited to settings compatible with single-agent reinforcement learning, which constrains their scalability and performance. Therefore, we propose QLLM, a multi-agent credit assignment algorithm based on the coder-evaluator framework, to address the aforementioned limitations.
Preliminaries
Decentralized Partially Observable Markov Decision Process
@@
In this work, we focus on a fully cooperative MARL setting that can be formalized as a decentralized partially observable Markov decision process (Dec-POMDP). A Dec-POMDP is typically defined by a tuple $G=\langle S,A,P,r,Z,O,n,\gamma\rangle$. Here, $n$ agents indexed by $i\in\{1,\dots,n\}$ operate in an environment with true underlying state $s\in S$. At each timestep, agent $i$ selects an action $a^{i}\in A$ based solely on its private observation $z_{i}\in Z$, which is sampled according to the observation function $O(s,i)$. The joint action of all agents is denoted by $\boldsymbol{a}\in\boldsymbol{A}$.
The environment evolves according to the transition probability function $P(s^{\prime}\mid s,\boldsymbol{a}):S\times\boldsymbol{A}\to[0,1]$, which specifies the likelihood of transitioning to state $s^{\prime}$ given the current state $s$ and joint action $\boldsymbol{a}$. All agents cooperate to maximize a shared reward signal defined by the function $r(s,\boldsymbol{a}):S\times\boldsymbol{A}\to\mathbb{R}$. The objective in this setting is to learn joint policies that maximize the expected cumulative discounted reward $\sum_{j=0}^{\infty}\gamma^{j}r_{t+j}$, where $\gamma\in[0,1)$ is a discount factor determining the weight of future rewards.
A key challenge in Dec-POMDPs lies in the credit assignment problem. This issue arises because all agents within the system share a common reward structure, making it inherently difficult to accurately attribute individual contributions to the collective outcome. Furthermore, the environment is perceived as non-stationary from the viewpoint of each agent. This non-stationarity poses significant challenges in achieving effective coordination and optimal decision-making, as agents must adapt to the changing strategies of others while managing incomplete state representations.
Figure 2: QLLM is designed within a coder-evaluator framework to autonomously generate high-quality, training-free credit assignment functions (TFCAFs). In each iteration, the coder LLM $M_{\text{coder}}$ produces $K$ candidate functions based on a task prompt and a role-specific instruction, referred to as the coder prompt. These candidate functions are initially assessed using environmental state information and local Q-values. If any runtime or syntax errors are identified, the faulty candidates are discarded and regenerated. Subsequently, the evaluator LLM $M_{\text{evaluator}}$ reviews the candidate functions, selects the most promising one based on its evaluator prompt This process is generated by LLMs through structured prompts without any need for human intervention.
Value Decomposition and IGM Principle
@@
Traditional value decomposition-based algorithms employ a mixing network to combine the local Q-value functions $Q_{i}$ of individual agents into a global Q-value function $Q_{\text{tot}}$. The mixing network serves as a differentiable function approximator that aggregates individual agent utilities while ensuring that the global Q-value adheres to the Individual-Global-Maximum (IGM) constraint. The IGM principle (Hong, Jin, and Tang 2022) requires that the joint action that maximizes the global action-value function $Q_{\text{tot}}$ must correspond to the combination of each agent’s individually optimal action. Formally, the IGM principle can be expressed as:
@@
where $\boldsymbol{\tau}\in T^{n}$ represents the joint action-observation histories of all agents. This equation implies that the maximization of the global Q-value function is aligned with the maximization of each agent’s local Q-value function, enabling decentralized execution while preserving global optimality. Building upon this, Qatten (Yang et al. 2020) provides a theoretical extension, showing that $Q_{tot}$ can be expressed as a linear combination of the individual $Q_{i}$ functions:
@@
Here, $w_{i,h}(s)$ denotes a coefficient that depends on the state $s$, $H$ is the number of attention heads. Theoretically, these coefficients can be derived from a higher-order functional expansion of $Q_{\text{tot}}$ with respect to $Q_{i}$, where the contribution of each term diminishes rapidly with increasing derivative order. This indicates that the influence of higher-order interaction terms on $Q_{\text{tot}}$ is sufficiently small to be ignored or effectively absorbed into a bias term $b$, allowing for a simplified model formulation and reduced computational complexity. Meanwhile, the theoretical insight in Equation (2) lays the groundwork for the subsequent development of TFCAF.
Method

### v5

# listing[Listing][List of Listings]

@@ Abstract
Credit assignment remains a fundamental challenge in multi-agent reinforcement learning (MARL) and is commonly addressed through value decomposition under the centralized training with decentralized execution (CTDE) paradigm. However, existing value decomposition methods typically rely on predefined mixing networks that require additional training, often leading to imprecise credit attribution and limited interpretability. We propose QLLM, a novel framework that leverages large language models (LLMs) to construct training-free credit assignment functions (TFCAFs), where the TFCAFs are nonlinear with respect to the global state and offer enhanced interpretability while introducing no extra learnable parameters. A coder-evaluator framework is employed to ensure the correctness and executability of the generated code. Extensive experiments on standard MARL benchmarks demonstrate that QLLM consistently outperforms baselines while requiring fewer learnable parameters. Furthermore, it demonstrates generalization across a broad set of value decomposition algorithms. Code is available at https://github.com/MaoMaoLYJ/pymarl-qllm.
1 Introduction
@@
Recent advances in cooperative multi-agent reinforcement learning (MARL) [28] have led to widespread applications in domains such as assembly lines [14], logistics distribution [42], autonomous driving [47], and swarm robotics [29]. In such scenarios, multiple agents learn to interact with the environment collaboratively to accomplish designated cooperative tasks. Since agents are typically trained using a shared team-level reward, accurately attributing individual contributions is critical for effective policy optimization [27]. This challenge is commonly referred to as the credit assignment problem, where inaccurate attribution can lead to suboptimal coordination behaviors, such as the emergence of lazy agents [22].
To address the credit assignment problem in multi-agent systems, a variety of approaches have been explored in the literature. Policy gradient approaches [1, 10, 6] address credit assignment through a centralized critic, where joint value or advantage functions are used to evaluate agents’ actions under shared rewards, and individual policies are updated based on agent-specific gradients derived from these centralized evaluations. In addtion, value decomposition methods [38] have emerged as a dominant paradigm under the centralized training with decentralized execution (CTDE) framework [23]. These methods construct a joint value function that is decomposed into individual utility functions, ensuring that the optimal joint policy can be recovered from local components while enabling scalable and coordinated training. In practice, value decomposition methods typically rely on neural-network-based mixing networks [35] to parameterize the credit assignment process. However, similar to other neural credit assignment approaches, they require additional training, incur substantial optimization overhead, and often lack semantic interpretability [21].
With the rapid advancement of large language models (LLMs) [13], they have demonstrated strong capabilities in natural language processing (NLP), including text understanding [2] and code generation [12]. Building on these capabilities, LLMs have increasingly been explored as general-purpose components in reinforcement learning (RL). In this context, LLMs can support RL in multiple ways, such as assisting reward design [25], constructing abstract world models [31], or directly serving as decision-making policies [18]. By injecting structured prior knowledge and high-level reasoning into the learning process, LLMs complement conventional data-driven optimization in complex environments.
@@
Figure 1: Comparison between the traditional value decomposition method using the mixing network (Left) and our proposed novel paradigm QLLM leveraging LLMs (Right).
Motivated by these developments, we investigate the use of LLMs to address the credit assignment problem in multi-agent systems. This paper introduces QLLM, a novel value decomposition framework that eliminates the need for a learned mixing network, as shown in Figure 1. To this end, a general coder-evaluator framework is proposed to support reliable LLM-based code generation. Within this framework, LLMs assume dual roles as both a code generator and a code evaluator [5], enabling the synthesis and validation of executable functions. Given task specifications and code generation instructions, the code generator produces a nonlinear Training-Free Credit Assignment Function (TFCAF), which serves as a direct substitute for the mixing network in value decomposition algorithms. Because TFCAF is generated directly by LLMs, it requires no additional training and offers improved semantic interpretability, while remaining robust to complex state representations. To further enhance reliability, a code verification mechanism is incorporated to ensure the correctness and executability of the generated code, thereby mitigating hallucination issues commonly observed in LLM-based code generation [26]. As a result, the proposed framework enables the construction of a dependable TFCAF without direct interaction with the environment, which can serve as a drop-in replacement for the mixing network and be readily integrated into a broad class of value decomposition methods.
The main contributions of this study are summarized as follows:
@@
    A coder-evaluator framework is adopted to enable the zero-shot construction of a reliable, training-free credit assignment function, termed TFCAF, via LLM-based code generation to the credit assignment problem in MARL.
@@
    A novel value decomposition framework, QLLM, is introduced by integrating TFCAF into MARL algorithms based on mixing network, allowing credit assignment without additional training or environment interaction.
@@
    Extensive experiments demonstrate that QLLM consistently outperforms existing credit assignment methods and provides improved semantic interpretability, particularly in complex state settings.
2 Related Work
2.1 Credit Assignment in MARL
@@
In recent years, the credit assignment problem has emerged as a fundamental challenge in multi-agent systems. Policy-based approaches [44, 45] were among the first to introduce credit assignment mechanisms. MADDPG [23] extends the DDPG [19] framework to multi-agent settings by incorporating CTDE. This approach utilizes an actor-critic architecture [15] with centralized critics, where the centralized critic has access to global information, enabling more effective learning in mixed cooperative-competitive scenarios. Similarly, COMA [6] adopts a centralized critic and introduces a counterfactual baseline to estimate agent-specific advantage functions. MAAC [10] further enhances policy-based methods by introducing an attention mechanism into the critic, leading to more precise credit assignment. However, most of these algorithms are not directly applicable to tasks with shared rewards.
@@
Another widely adopted paradigm is the value decomposition approach [35, 41], which explicitly decomposes the globally shared return. As one of the earliest methods in this category, VDN [38] models the global value function as a linear summation of individual value functions. While simple and computationally efficient, VDN struggles to capture complex cooperative relationships due to its linear decomposition. QMIX [35] improves upon VDN by using a monotonic mixing network to non-linearly combine individual value functions, allowing for more expressive representations. However, the assumption of monotonicity between individual and global values may restrict its flexibility in highly complex environments. QTRAN [37] and QPLEX [40] mitigate this limitation by relaxing the monotonicity constraint, enabling the model to learn richer representations, albeit at the cost of increased computational complexity. RIIT [9] goes further by modeling implicit influence relationships between agents, enabling better credit assignment in tasks with long-term dependencies. Nevertheless, RIIT ultimately retains a mixing network structure analogous to that used in QMIX, in order to establish the mapping from individual agent Q-values to the global Q-value. While the mixing network plays a central role in implementing credit assignment, it often suffers from limited interpretability and slow convergence during training. As a result, despite the success of value decomposition methods, achieving efficient, precise, and interpretable credit assignment in multi-agent systems remains an open challenge.
2.2 LLM-Enhanced RL
Leveraging their large-scale pretraining and strong generalization capabilities, LLMs have been increasingly integrated into RL to enhance performance in multitask learning, sample efficiency, and high-level task planning [46, 17]. Recent studies have explored the use of LLMs to generate structured decision policies or programmatic representations from task descriptions, enabling more interpretable and adaptable control strategies [4, 18]. In addition, LLMs can provide natural language rationales for agent behaviors, improving the transparency and explainability of decision-making processes [3, 32]. LLMs have been investigated as tools for shaping reward signals, where automatically generated reward functions or feedback mechanisms help guide agents toward more effective behaviors [25, 24]. Collectively, these works highlight the potential of LLMs as flexible components for enhancing learning signals and structural priors in RL.
@@ 2.2 LLM-Enhanced RL
Recent work has begun to explore the use of LLMs for temporal and agent-level credit assignment in MARL. Qu et al. [33] proposed latent rewards to capture cooperative relationships, effectively addressing credit assignment under sparse rewards and outperforming existing baselines. Meanwhile, Lin et al. [20] employed LLMs to learn potential-based reward functions across multiple queries, allowing for more accurate estimation of individual agent contributions. Despite their effectiveness, these methods primarily focus on reward decomposition and remain limited to settings compatible with single-agent reinforcement learning formulations, which restricts their scalability and applicability to general MARL algorithms. To overcome these limitations, QLLM is introduced as a multi-agent credit assignment framework that uses LLM-based code generation to construct training-free credit assignment functions compatible with value decomposition methods.
3 Preliminaries
3.1 Dec-POMDP
@@
In this work, we focus on a fully cooperative MARL setting formalized as a decentralized partially observable Markov decision process (Dec-POMDP[27]. A Dec-POMDP is typically defined by a tuple $G=\langle S,A,P,r,Z,O,n,\gamma\rangle$. In this setting, $n$ agents indexed by $i\in\{1,\dots,n\}$ operate in an environment with a true underlying state $s\in S$. At each timestep, agent $i$ selects an action $a^{i}\in A$ based solely on its private observation $z_{i}\in Z$, which is sampled according to the observation function $O(s,i)$. The joint action of all agents is denoted by $\boldsymbol{a}\in\boldsymbol{A}$. The environment evolves according to the transition probability function $P(s^{\prime}\mid s,\boldsymbol{a}):S\times\boldsymbol{A}\to[0,1]$, which specifies the probability of transitioning to state $s^{\prime}$ given the current state $s$ and joint action $\boldsymbol{a}$. All agents cooperate to maximize a shared reward signal defined by the reward function $r(s,\boldsymbol{a}):S\times\boldsymbol{A}\to\mathbb{R}$. The objective in this setting is to learn joint policies that maximize the expected cumulative discounted reward $\sum_{j=0}^{\infty}\gamma^{j}r_{t+j}$, where $\gamma\in[0,1)$ is a discount factor.
Dec-POMDPs pose inherent challenges due to partial observability and non-stationarity. Beyond these difficulties, a central challenge in Dec-POMDPs is the credit assignment problem, which is the primary focus of this work. Because agents typically optimize a shared team reward, accurately attributing individual contributions to the collective outcome is non-trivial and plays a critical role in effective coordination and learning.
3.2 Value Decomposition and IGM Principle
@@
Traditional value decomposition methods employ a mixing network to aggregate the local action-value functions $Q_{i}$ of individual agents into a global action-value function $Q_{\text{tot}}$ following the Individual-Global-Maximum (IGM) principle [35],whereby the joint action that maximizes $Q_{\text{tot}}$ must correspond to the combination of each agent’s individually optimal action. Formally, the IGM principle can be expressed as:
@@
where $\boldsymbol{\tau}=(\tau^{1},\dots,\tau^{n})$ represents the joint action-observation histories of all agents. This equation implies that the maximization of the global Q-value function is aligned with the maximization of each agent’s local Q-value function, enabling decentralized execution while preserving global optimality. Building upon this principle, Qatten [43] provides a theoretical extension, showing that $Q_{\text{tot}}$ can be expressed as a linear combination of the individual $Q_{i}$ functions:
@@
In this formulation, $w_{i,h}(s)$ denotes a coefficient that depends on the state $s$, and $H$ is the number of attention heads. Theoretically, these coefficients can be derived from a higher-order functional expansion of $Q_{\text{tot}}$ with respect to $Q_{i}$, where the contribution of each term diminishes rapidly with increasing derivative order. This indicates that the influence of higher-order interaction terms on $Q_{\text{tot}}$ is sufficiently small to be ignored or effectively absorbed into a bias term $b$, allowing for a simplified model formulation and reduced computational complexity.
4 Methodology